OpenStudy (anonymous):

PLEASE HELP =) Write x2 - 6x + 7 = 0 in the form (x - a)2 = b, where a and b are integers.

OpenStudy (anonymous):

(x - 4)2 = 3 (x - 3)2 = 2 (x - 2)2 = 1 (x - 1)2 = 4

rishavraj (rishavraj):

-6x = 2*(-3)*x

rishavraj (rishavraj):

proceed

OpenStudy (anonymous):

Uhhhhh sorry =)

rishavraj (rishavraj):

for wth????o_O @CountryGurl15

OpenStudy (anonymous):

I'm sorry you said proceed and idk what to do next =)

rishavraj (rishavraj):

(x - a)^2 = x^2 + a^2 - 2ax see the term having negative sign is formed by multiplying all given terms with 2....

OpenStudy (anonymous):

Oh okay =)

rishavraj (rishavraj):

here u r having -6x which mean its 2*(-3)x

rishavraj (rishavraj):

so u got "a" a = 3 proceed

OpenStudy (anonymous):

Ohhh okay so we get -5x

rishavraj (rishavraj):

wht u mean??? in (x - a)^2 a = 3.......

OpenStudy (anonymous):

I'm sorry if that's wrong, i'm trying to get this i really am =)

OpenStudy (anonymous):

oh so it's (x-3)^2 ?

rishavraj (rishavraj):

nah... u r close..see in question given equation is x2 - 6x + 7 = 0

OpenStudy (anonymous):

(x-3)^2 = 2?

OpenStudy (anonymous):

Yes. . . . .

rishavraj (rishavraj):

@CountryGurl15 bingo u r right...:))

OpenStudy (anonymous):

Thanks so much! =)

rishavraj (rishavraj):

u r most welcome......:)